An engineer wishes to determine the width of a particular electronic component. If she knows that the standard deviation is 1.3 mm, how many of these components should she consider to be 80% sure of knowing the mean will be within ± 0.1 mm?
Answer #1

Solution :

Given that,

standard deviation = =1.3

Margin of error = E = +/-0.1

At 80% confidence level the z is ,

= 1 - 80% = 1 - 0.80 = 0.20

/ 2 =0.10

Z/2 = Z0.10= 1.28 Using z table    )

sample size = n = [Z/2* / E] 2

n = ( 1.28* 1.3 / 0.1 )2

n =276.8

Sample size = n =277